OpenStudy (pokemon23):

Find two consecutive even integers such that the square of the smaller is 10 more than the larger.

OpenStudy (pokemon23):

I'm a bit confused in solving this

OpenStudy (mani_jha):

let the nos be x and x+2 x^2-(x+2)=10

OpenStudy (mani_jha):

you might have thought it was : x^2-(x+2)^2. isnt it?

OpenStudy (pokemon23):

ya

OpenStudy (pokemon23):

=10

OpenStudy (mani_jha):

ya, that's a grammatical problem in the question.

OpenStudy (mani_jha):

hey where r u from?

OpenStudy (pokemon23):

I'm from NY my teacher just can't write out word problem

OpenStudy (pokemon23):

I still don't get how you it...

Directrix (directrix):

Something is wrong with my work. Anybody know? Let x and x+2 be consecutive even integers given that x is even. x^2 = 10(x+2) x^2 = 10x + 20 x^2 – 10 x – 20 = 0

OpenStudy (mani_jha):

Two even consecutive integers, they are like 2,4 or 4,6 or 8,10 or just x, x+2(where x is an even number). Is this clear? The smaller number is x and larger is x+2. The problem says that the square of the smaller number(x^2) is 10 more than the larger number(x+2). x^2=10+(x+2) or x^2-(x+2)=10 Please tell me if you didnt understand. @directrix, The square of the smaller no is 1O more than the larger number, not 10 times the larger number.

OpenStudy (pokemon23):

I understand the first statement

Directrix (directrix):

Ten MORE than, not ten TIMES. @Mani --> I read incorrectly. Thanks for telling me. Now, I will re-work the problem.

Directrix (directrix):

So, 4 and 6 would be one pair. Then, I got -3 but it is odd and is still odd when I add 2 and get -1. Therefore, x = -3 is an extraneous root for this scenario. Yes?

OpenStudy (mani_jha):

The answer is 4,6. See that the square of 4 is 16, which is 10 more than 6. In algebra, you have got to treat values like x and x+2 just as normal numbers. So, if x is a no, its square is x^2. And if it is 10 more than the lager no. x^2=10+(x+2){compare this with 16=10+6} @directrix, x=-3 is a solution for the equation but not of this problem, because the problem asks for even integers. Is that the meaning of 'extraneous'?

OpenStudy (pokemon23):

Never knew things could be become complicated than it seems

Directrix (directrix):

I thought so but then I read this. Extraneous Solution Spurious Solution A solution of a simplified version of an equation that does not satisfy the original equation. Watch out for extraneous solutions when solving equations with a variable in the denominator of a rational expression, with a variable in the argument of a logarithm, or a variable as the radicand in an nth root when n is an even number. http://www.mathwords.com/e/extraneous_solution.htm

OpenStudy (mani_jha):

@pokemon23 , i am sorry that i am unable to explain this to you. But can you tell me what is it about the problem that you dont get?

OpenStudy (mani_jha):

I mean, the exact part where u dont understand

Directrix (directrix):

While the Poke is thinking, I'll post a problem about sequences. Maybe you will work on it. Just a sec.

OpenStudy (pokemon23):

umm Let after you got x^2=10x(x+2) how did you get x= to 4 and 6

OpenStudy (mani_jha):

Oh, sorry I skipped a few steps. x^2=10+(x+2) x^2-x-2=10 x^2-x-12=0 x^2-4x+3x-12=0 x(x-4)+3(x-4)=0 (x+3)(x-4)=0 So, you get either x=-3 or x=4. But the question asks for even integers, so we get only x=4. And so the even integer consecutive to it is x+2=6. Now, is it ok?

OpenStudy (pokemon23):

ya ok because I was wondering like how did he do this thanks for the help and direct as well

Directrix (directrix):

@ Mani, I get it, too.. Thanks. Will you work on the problem I just posted?